# RLHF and IIA: Perverse Incentives

### 02 Dec 2023 v1

@@ 1 Introduction
Modern generative AIs ingest trillions of data bytes from the World Wide Web to produce a large pretrained model. Trained to imitate what is observed, this model represents an agglomeration of behaviors, some of which are more or less desirable to mimic. Further training through human interaction, even on fewer than a hundred thousand bits of data, has proven to greatly enhance usefulness and safety, enabling the remarkable AIs we have today. This process of reinforcement learning from human feedback (RLHF) steers AIs toward the more desirable among behaviors observed during pretraining.
While AIs now routinely generate drawings, music, speech, and computer code, the text-based chatbot remains an emblematic artifact. To produce a chatbot, starting with a pretrained language model, a prototypical approach to RLHF [4, 13, 9] progresses through several steps. First, queries, each taking the form of a prompt and a pair of alternative responses, are presented to human annotators who each identify their favorite among a pair. The annotated data is then used to train a reward model to score any response to a given prompt. Finally, the language model is tuned to align its responses toward those that earn high reward.
The tremendous impact of RLHF in generative AI has sparked a flurry of research that aims to understand and improve the process. Some propose alternative algorithms [10, 14, 16, 5]. Others consider alternative query formats [17, 12, 15, 10] for which annotators, rather than comparing only a pair of responses, are asked to choose from a longer list, or to rank-order. Feedback can also be garnered from interactions with humans in their regular use of online chatbots. RLHF research is continually growing in importance with the volume of human feedback data.
With all the effort and resources directed at RLHF, it is worth asking whether current algorithms rest on firm foundations. Maybe not, as these algorithms are based on models that assume independence of irrelevant alternatives (IIA), which does not cohere with human preferences for text content. Interestingly, this flaw is not pronounced when using the most common approach of fitting a reward model to pairwise comparison data, then tuning a language model to optimize reward. However, working with alternative query formats or algorithmic innovations can induce egregious behavior.
To illustrate such egregious behavior, let us consider a simplified language model that only generates messages explaining a particular subject. Each message can be categorized as concise or verbose. Suppose that $60\%$ of the population prefers concise over verbose explanations, but individuals are indifferent between explanations within each category. In a representative query, a set of alternatives $\mathcal{Y}$ is presented to an annotator, who selects an element $y\in\mathcal{Y}$. Suppose that elements of $\mathcal{Y}$ are drawn randomly from corpi of concise and verbose messages, with samples among each category drawn uniformly. If both categories are represented in $\mathcal{Y}$, a random annotator chooses a concise message with probability $0.6$, and otherwise, a verbose one. Conditioned on the category, the annotator’s choice is uniformly distributed across the intersection of $\mathcal{Y}$ and that category.
Figure 1 plots results from applying existing RLHF algorithms to a didactic model of this kind, details of which are provided in Section 5.1. The graph on the left was produced by a standard RLHF algorithm [4, 13, 9] that carries out reward learning followed by policy optimization, which we refer to as RLPO. With pairwise comparisons, where each query includes only two alternatives, as the number of queries grows, resulting policies always produce concise messages. This is an intended outcome, as concise messages are preferred by the majority, and is indicated by convergence of the blue plot to probability one. However, when queries include larger numbers of alternatives, policies produced by RLPO consistently generate verbose messages. This failure mode is reflected by the other plots on the graph, which converge to probability zero.
Figure 1: Existing RLHF algorithms can generate egregious behavior.
The preceding discussion indicates that presenting more than two messages to each annotator can induce egregious behavior. The graph on the right illustrates how, even with pairwise comparisons, innovating on RLHF algorithms can result in undesirable outcomes. These results were produced by inclusive learning (IL), a new algorithm that aims to match human choice probabilities [2, 14]. This is in contrast to RLPO, which targets the most common choice. In our didactic example, if the algorithms work as intended, IL should generate concise messages $60\%$ of the time, as indicated by the horizontal dotted line, while RLPO should do so all the time. The plots indicate that, for any fixed corpus of concise messages, as the corpus of verbose messages grows, policies produced by IL consistently generate verbose messages. This occurs irrespective of the number of alternatives $|\mathcal{Y}|$ presented to each annotator.
The egregious behaviors appearing with queries that extend beyond pairwise comparisons or algorithmic innovations like IL stem from the IIA assumption inherent to models underlying RLHF algorithms. The remainder of the paper explains issues and methods more deeply, with a goal to enhance understanding of flaws in current algorithms. We develop this understanding through interpreting simulation results of the sort we have presented as well as theoretical results that corroborate the qualitative behavior. These theoretical results establish that such behavior generalizes beyond the particular simulated instances. We leave for future work the design of algorithms that alleviate flaws we identify here.
@@ 2.2 Messages
There is a set $\mathcal{M}\subseteq\mathcal{A}^{+}$ of token sequences that we refer to as messages. We interpret $\mathcal{M}$ as the set of complete individual statements. For example, in a text messaging application, a message might terminate with either the ASCII enter character or upon reaching a maximum allowed message length.
A language model generates a random message by sampling sequentially according to $X_{t+1}\sim\pi(\cdot|(X_{1},\ldots,X_{t}))$, terminating at the first index $T$ for which $(X_{1},\ldots,X_{T})\in\mathcal{M}$. For any token sequence $x=(x_{1},\ldots,x_{T})\in\mathcal{A}^{+}$, we use the notation $x_{t_{1}:t_{2}}$ to denote the subsequence $(x_{t_{1}},\ldots,x_{t_{2}})$. Hence, we can use the abbreviated form $X_{t+1}\sim\pi(\cdot|X_{1:t})$. We assume that, for any language model $\pi$, any token sequence generated sequentially in this manner terminates with probability one. Through this procedure, a language model $\pi$ samples messages from a distribution $P_{\pi}$, defined by
@@ 2.5 What About Prompting?
The RLHF literature typically consider generation of a message in response to previous messages, referred to in aggregate as a prompt. In that context, each preference datum expresses a choice between responses to a prompt. We omit prompting from our formulation and analysis, because that would only complicate the discussion without contributing to insight. The examples, algorithms, and results we present can easily be extended to treat prompting.
@@ 3.2 Reward Functions
A reward function $r$ expresses individual preferences between elements of a message set $\mathcal{M}$. In particular, an individual of type $z\in\mathcal{Z}$ prefers a message $x$ to $x^{\prime}$ if and only if $r(x|z)\geq r(x^{\prime}|z)$.
To illustrate with an example, two elements $z,z^{\prime}\in\mathcal{Z}$ could express detailed characteristics of particular individuals, one of whom resides in China and the other in the United States. These individuals have different tastes and needs – a message $x$ that is helpful or interesting to one may not be to the other. The distinct preferences are expressed by differing reward assignments $r(x|z)\neq r(x|z^{\prime})$.
@@ 3.4 Example 1: Logit Models
Conversely, only the choice model $(\mathcal{Z},p,r)$ that we have specified or one with a reward function that differs from $r$ by a constant produces these choice probabilities [8].
As an example, consider a language model $\pi$ that recommends recipes for dinner. For each recipe $y$, $\overline{r}(y)$ indicates its nominal desirability – no individual attributes any less reward to $y$. But for a specific individual of type $z$, a positive perturbation $z(y)$ is added to reflect their idiosyncratic preferences. Presented with a set $\mathcal{Y}$ of recipes, the individual selects the recipe $y\in\mathcal{Y}$ that maximizes $\overline{r}(y)+z(y)$. In this way, $z$ induces variation in how different individuals make their selection. The Gumbel distribution enjoys elegant mathematical structure: if, for each recipe $y$, $z(y)$ is an independent sample from the standard Gumbel then the probability that a random individual selects $y$ over the remaining recipes in $\mathcal{Y}$ is given by (3).
@@ 3.5 Example 2: Soft Choice Models
instead of (2).
A soft choice model $(\mathcal{Z},p,r)$ can be expressed as a hard choice model $(\mathcal{Z}^{\prime},p^{\prime},r^{\prime})$. By this we mean that choice probabilities $\mathbb{P}(Y=y|\mathcal{Y})$ implied by the two models are identical. The type set $\mathcal{Z}^{\prime}$ is the product set $\mathcal{Z}\times\Re_{+}^{\mathcal{M}}$. The type of a random individual is a pair $Z^{\prime}=(Z,W)$. The first item $Z$ is sampled from $p$. The second item comprises of an independent standard Gumbel sample $W(x)$ for each $x\in\mathcal{M}$. The reward function is given by $r^{\prime}(x|z^{\prime})=r(x|z)+w(x)$. It is easy to see that this hard choice model $(\mathcal{Z}^{\prime},p^{\prime},r^{\prime})$ produces the same choice probabilities (4) as the soft choice model $(\mathcal{Z},p,r)$.
Note that a logit model can be expressed as a soft choice model. Consider a logit model expressed as a hard choice model $(\mathcal{Z},p,r)$. Define a soft choice model $(\mathcal{Z}^{\prime},p^{\prime},r^{\prime})$ with a single type $\mathcal{Z}=\{0\}$, a trivial type distribution for which $p(0)=1$, and a reward function $r^{\prime}(x|0)=r(x|0)$. This soft choice model implies choice probabilties
@@ 3.6 Example 3: Dichotomy Models
Note that reward depends on the message $x$ only through its membership in $\mathcal{M}_{1}$ or $\mathcal{M}_{2}$. In other words, each individual is indifferent between any two messages of the same type.
The type set $\mathcal{Z}$ and reward function $r$ we have defined, together with any type distribution $p$, constructs a choice model $(\mathcal{Z},p,r)$. According to this choice model, when presented with alternatives $\mathcal{Y}$, the choice $Y$ made by an individual of type $Z$ is sampled uniformly from $\mathcal{Y}\cap\mathcal{M}_{Z}$. Hence,
@@
where $Z$ is sampled from $p$.
A hypothetical sort of homophily serves to illustrate a context where this sort of model may be relevant. Suppose a language model $\pi$ generates political commentary, with $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ conveying conservative versus liberal views. In this hypothetical world, there are two types of individuals – conservatives and liberals – labeled $1$ and $2$. Each prefers commentary aligned with their predisposition. There is no reward for opposing views.
@@ 3.7 Independence of Irrelevant Alternatives
Consider three messages =$\mathcal{M}=\{\mathrm{dog},\mathrm{cat},\mathrm{feline}\}$ that each guess an individual’s favorite pet. Consider a logit choice model, as described in Section 3.4, that assigns equal base rewards $r(\mathrm{dog})=r(\mathrm{cat})=r(\mathrm{feline})$ to indicate that a random individual is as likely to prefer dogs or cats. Presented with two alternatives $\mathcal{Y}=\{\mathrm{dog},\mathrm{cat}\}$, the logit model produces choice probabilities $\mathbb{P}(Y=\mathrm{dog}|\mathcal{Y})=\mathbb{P}(Y=\mathrm{cat}|\mathcal{Y})=1/2$. On the other hand, with three alternatives $\mathcal{Y}=\{\mathrm{dog},\mathrm{cat},\mathrm{feline}\}$, $\mathbb{P}(Y=\mathrm{dog}|\mathcal{Y})=\mathbb{P}(Y=\mathrm{cat}|\mathcal{Y})=\mathbb{P}(Y=\mathrm{feline}|\mathcal{Y})=1/3$. This of course makes no sense, because feline is synonymous to cat, so presenting it as an alternative ought not reduce the probability that an individual prefers dogs.
The aforementioned implausible implication of a logit model derives from its assumption of independence of irrelevant alternatives (IIA). While this notion was formalized in earlier work [1], the treatment of [6] best suits our usage. In that treatment, IIA indicates that the ratio of any two choice probabilities does not vary with the set of alternatives. More formally, for all messages $x,x^{\prime}\in\mathcal{M}$ and sets $\mathcal{X}$ and $\mathcal{X}^{\prime}$ such that $x,x^{\prime}\in\mathcal{X}\cap\mathcal{X}^{\prime}$,
@@ 3.8 IIA for Language
Human preferences for text messages do not satisfy IIA. This is because, for any given message, others can be virtually equivalent. Our preceding example with three messages – dog, cat, and feline – illustrates this. The introduction of feline as an alternative should not reduce the probability that an individual chooses dog.
In spite of this, current RLHF algorithms are based on models that assume IIA. In particular, preference estimates used in RLHF satisfy IIA even though human preferences expressed through data do not. These estimates can give rise to perverse incentives to generate undesirable messages. We will discuss in Section 5 potentially egregious consequences. But first, we review current RLHF algorithms.
@@ 4.1 Reward Learning Followed by Policy Optimization
The reward function $r_{\psi}$ could, for example, be linear in a vector message embedding $\phi(x)$, taking the form $\widehat{r}_{\psi}(x)=\psi^{\top}\phi(x)$. Alternatively, $\widehat{r}_{\psi}$ could be a neural network with parameters $\psi$.
For any $r:\mathcal{M}\rightarrow\Re$ and language model $\pi$, we define a policy loss function
@@
for some scalar hyperparameter $\beta>0$. The first term of this loss function expresses expected reward while the second regularizes toward the base language model.
A fine-tuned language model is obtained via two steps. The first minimizes $\mathcal{L}_{\mathrm{reward}}$ to produce a reward function $\widehat{r}_{\widehat{\psi}}$. This is followed by a policy optimization step, which, given a language model $\widehat{\pi}_{\theta}$ parameterized by a vector $\theta$, minimizes $\mathcal{L}_{\mathrm{policy}}(\cdot|\widehat{r}_{\widehat{\psi}})$ to produce the fine-tuned language model $\widehat{\pi}_{\widehat{\theta}}$. This minimization is typically carried out via stochastic gradient descent starting with $\theta$ initialized to match the base policy $\widehat{\pi}_{\theta}=\overline{\pi}$.

### 21 Dec 2023 v2

@@ 1 Introduction
Modern generative AIs ingest trillions of data bytes from the World Wide Web to produce a large pretrained model. Trained to imitate what is observed, this model represents an agglomeration of behaviors, some of which are more or less desirable to mimic. Further training through human interaction, even on fewer than a hundred thousand bits of data, has proven to greatly enhance usefulness and safety, enabling the remarkable AIs we have today. This process of reinforcement learning from human feedback (RLHF) steers AIs toward the more desirable among behaviors observed during pretraining. While AIs now routinely generate drawings, music, speech, and computer code, the text-based chatbot remains an emblematic artifact. To produce a chatbot, starting with a pretrained language model, a prototypical approach to RLHF [4, 15, 11] progresses through several steps. First, queries, each taking the form of a prompt and a pair of alternative responses, are presented to human annotators who each identify their favorite among a pair. The annotated data is then used to train a reward model to score any response to a given prompt. Finally, the language model is tuned to align its responses toward those that earn high reward. The tremendous impact of RLHF in generative AI has sparked a flurry of research that aims to understand and improve the process. Some propose alternative algorithms [12, 16, 18, 6]. Others consider alternative query formats [19, 14, 17, 12] for which annotators, rather than comparing only a pair of responses, are asked to choose from a longer list, or to rank-order. Feedback can also be garnered from interactions with humans in their regular use of online chatbots. RLHF research is continually growing in importance with the volume of human feedback data. With all the effort and resources directed at RLHF, it is worth asking whether current algorithms rest on firm foundations. Maybe not, as these algorithms are based on models that assume independence of irrelevant alternatives (IIA), which does not cohere with human preferences for text content. Interestingly, this flaw is not pronounced when using the most common approach of fitting a reward model to pairwise comparison data, then tuning a language model to optimize reward. However, working with alternative query formats or algorithmic innovations can induce egregious behavior. A simple experiment we will present in Section 6 illustrates our point. This experiment applies a standard reward learning algorithm [4, 15, 11]. We first consider learning from queries that are each comprised of the prompt
prompt Did Oppenheimer win a Nobel Prize?
and a pair of responses, one generated by GPT-3.5 and the other by GPT-4. The former generally produces more concise and the latter more informative responses. For example, here are representative responses:
GPT-3.5 response No, Oppenheimer did not win the Nobel Prize. GPT-4 response No, Robert Oppenheimer, often called the “father of the atomic bomb” for his role in the Manhattan Project, did not win a Nobel Prize.
If a large majority of annotators prefer responses generated by GPT-4, the learned reward function correctly assigns higher scores to GPT-4 responses in independent test data. A variation in which training queries include four rather than two responses reveals egregious behavior induced by the IIA assumption. In particular, suppose that each query includes one response generated by GPT-3.5 and three by GPT-4. Then, even if a large majority of annotators prefer responses generated by GPT-4, the learned reward function erroneously assigns higher scores to GPT-3.5 responses in test data. Aside from innocuous changes to the query format, innovating on RLHF algorithms can also result in undesirable behavior. To offer an example, we will consider inclusive learning (IL), an approach that aims to produce a language model that samples responses according to human choice probabilities [2, 16]. This is in contrast to the standard approach to RLHF, which targets responses with the highest human choice probabilities. In Section 4, we will demonstrate that, even with two responses per query, the IIA assumption exacts egregious behavior from IL. The remainder of the paper explains issues and methods more deeply, with a goal to enhance understanding of flaws in current algorithms. We develop this understanding through interpreting simulations of didactic models, an empirical study of data produced by GPT-3.5 and GPT-4, and theoretical results that corroborate the observed behavior. These theoretical results establish that such behavior generalizes beyond specific instances. We leave for future work the design of algorithms that alleviate flaws we identify here.
@@ 2.2 Messages
There is a set $\mathcal{M}\subseteq\mathcal{A}^{+}$ of token sequences that we refer to as messages. We interpret $\mathcal{M}$ as the set of complete individual statements. For example, in a text messaging application, a message might terminate with either the ASCII enter character or upon reaching a maximum allowed message length. A language model generates a random message by sampling sequentially according to $X_{t+1}\sim\pi(\cdot|(X_{1},\ldots,X_{t}))$, terminating at the first index $T$ for which $(X_{1},\ldots,X_{T})\in\mathcal{M}$. For any token sequence $x=(x_{1},\ldots,x_{T})\in\mathcal{A}^{+}$, we use the notation $x_{t_{1}:t_{2}}$ to denote the subsequence $(x_{t_{1}},\ldots,x_{t_{2}})$. Hence, we can use the abbreviated form $X_{t+1}\sim\pi(\cdot|X_{1:t})$. We assume that, for any language model $\pi$, any token sequence generated sequentially in this manner terminates with probability one. Through this procedure, a language model $\pi$ samples messages from a distribution $P_{\pi}$, defined by
@@ 2.5 What About Prompting?
Language models typically generate a message in response to a prompt made up of previous messages. In that context, each preference datum expresses a choice between responses to a prompt. Except for our empirical study of Section 6, we omit prompting from our formulation and analysis, because that would only complicate the discussion without contributing to insight. Examples, algorithms, and results we present can easily be extended to treat prompting.
@@ 3.2 Reward Functions
A reward function $r$ expresses individual preferences between elements of a message set $\mathcal{M}$. In particular, an individual of type $z\in\mathcal{Z}$ prefers a message $x$ to $x^{\prime}$ if and only if $r(x|z)\geq r(x^{\prime}|z)$. To illustrate with an example, two elements $z,z^{\prime}\in\mathcal{Z}$ could express detailed characteristics of particular individuals, one of whom resides in China and the other in the United States. These individuals have different tastes and needs – a message $x$ that is helpful or interesting to one may not be to the other. The distinct preferences are expressed by differing reward assignments $r(x|z)\neq r(x|z^{\prime})$.
@@ 3.4 Example 1: Logit Models
Conversely, only the choice model $(\mathcal{Z},p,r)$ that we have specified or one with a reward function that differs from $r$ by a constant produces these choice probabilities [9]. As an example, consider a language model $\pi$ that recommends recipes for dinner. For each recipe $y$, $\overline{r}(y)$ indicates its nominal desirability – no individual attributes any less reward to $y$. But for a specific individual of type $z$, a positive perturbation $z(y)$ is added to reflect their idiosyncratic preferences. Presented with a set $\mathcal{Y}$ of recipes, the individual selects the recipe $y\in\mathcal{Y}$ that maximizes $\overline{r}(y)+z(y)$. In this way, $z$ induces variation in how different individuals make their selection. The Gumbel distribution enjoys elegant mathematical structure: if, for each recipe $y$, $z(y)$ is an independent sample from the standard Gumbel then the probability that a random individual selects $y$ over the remaining recipes in $\mathcal{Y}$ is given by (3).
@@ 3.5 Example 2: Soft Choice Models
instead of (2). A soft choice model $(\mathcal{Z},p,r)$ can be expressed as a hard choice model $(\mathcal{Z}^{\prime},p^{\prime},r^{\prime})$. By this we mean that choice probabilities $\mathbb{P}(Y=y|\mathcal{Y})$ implied by the two models are identical. The type set $\mathcal{Z}^{\prime}$ is the product set $\mathcal{Z}\times\Re_{+}^{\mathcal{M}}$. The type of a random individual is a pair $Z^{\prime}=(Z,W)$. The first item $Z$ is sampled from $p$. The second item comprises of an independent standard Gumbel sample $W(x)$ for each $x\in\mathcal{M}$. The reward function is given by $r^{\prime}(x|z^{\prime})=r(x|z)+w(x)$. It is easy to see that this hard choice model $(\mathcal{Z}^{\prime},p^{\prime},r^{\prime})$ produces the same choice probabilities (4) as the soft choice model $(\mathcal{Z},p,r)$. Note that a logit model can be expressed as a soft choice model. Consider a logit model expressed as a hard choice model $(\mathcal{Z},p,r)$. Define a soft choice model $(\mathcal{Z}^{\prime},p^{\prime},r^{\prime})$ with a single type $\mathcal{Z}=\{0\}$, a trivial type distribution for which $p(0)=1$, and a reward function $r^{\prime}(x|0)=r(x|0)$. This soft choice model implies choice probabilties
@@ 3.6 Example 3: Dichotomy Models
Note that reward depends on the message $x$ only through its membership in $\mathcal{M}_{1}$ or $\mathcal{M}_{2}$. In other words, each individual is indifferent between any two messages of the same type. The type set $\mathcal{Z}$ and reward function $r$ we have defined, together with any type distribution $p$, constructs a choice model $(\mathcal{Z},p,r)$. According to this choice model, when presented with alternatives $\mathcal{Y}$, the choice $Y$ made by an individual of type $Z$ is sampled uniformly from $\mathcal{Y}\cap\mathcal{M}_{Z}$. Hence,
@@
where $Z$ is sampled from $p$. A hypothetical sort of homophily serves to illustrate a context where this sort of model may be relevant. Suppose a language model $\pi$ generates political commentary, with $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ conveying conservative versus liberal views. In this hypothetical world, there are two types of individuals – conservatives and liberals – labeled $1$ and $2$. Each prefers commentary aligned with their predisposition. There is no reward for opposing views.
@@ 3.7 Independence of Irrelevant Alternatives
Consider three messages =$\mathcal{M}=\{\mathrm{dog},\mathrm{cat},\mathrm{feline}\}$ that each guess an individual’s favorite pet. Consider a logit choice model, as described in Section 3.4, that assigns equal base rewards $r(\mathrm{dog})=r(\mathrm{cat})=r(\mathrm{feline})$ to indicate that a random individual is as likely to prefer dogs or cats. Presented with two alternatives $\mathcal{Y}=\{\mathrm{dog},\mathrm{cat}\}$, the logit model produces choice probabilities $\mathbb{P}(Y=\mathrm{dog}|\mathcal{Y})=\mathbb{P}(Y=\mathrm{cat}|\mathcal{Y})=1/2$. On the other hand, with three alternatives $\mathcal{Y}=\{\mathrm{dog},\mathrm{cat},\mathrm{feline}\}$, $\mathbb{P}(Y=\mathrm{dog}|\mathcal{Y})=\mathbb{P}(Y=\mathrm{cat}|\mathcal{Y})=\mathbb{P}(Y=\mathrm{feline}|\mathcal{Y})=1/3$. This of course makes no sense, because feline is synonymous to cat, so presenting it as an alternative ought not reduce the probability that an individual prefers dogs. The aforementioned implausible implication of a logit model derives from its assumption of independence of irrelevant alternatives (IIA). While this notion was formalized in earlier work [1], the treatment of [7] best suits our usage. In that treatment, IIA indicates that the ratio of any two choice probabilities does not vary with the set of alternatives. More formally, for all messages $x,x^{\prime}\in\mathcal{M}$ and sets $\mathcal{X}$ and $\mathcal{X}^{\prime}$ such that $x,x^{\prime}\in\mathcal{X}\cap\mathcal{X}^{\prime}$,
@@ 3.8 IIA for Language
Human preferences for text messages do not satisfy IIA. This is because, for any given message, others can be virtually equivalent. Our preceding example with three messages – dog, cat, and feline – illustrates this. The introduction of feline as an alternative should not reduce the probability that an individual chooses dog. In spite of this, current RLHF algorithms are based on models that assume IIA. In particular, preference estimates used in RLHF satisfy IIA even though human preferences expressed through data do not. These estimates can give rise to perverse incentives to generate undesirable messages. We will discuss in Section 5 potentially egregious consequences. But first, we review current RLHF algorithms.
@@ 4.1 Reward Learning Followed by Policy Optimization
The reward function $r_{\psi}$ could, for example, be linear in a vector message embedding $\phi(x)$, taking the form $\widehat{r}_{\psi}(x)=\psi^{\top}\phi(x)$. Alternatively, $\widehat{r}_{\psi}$ could be a neural network with parameters $\psi$. For any $r:\mathcal{M}\rightarrow\Re$ and language model $\pi$, we define a policy loss function
@@
for some scalar hyperparameter $\beta>0$. The first term of this loss function expresses expected reward while the second regularizes toward the base language model. A fine-tuned language model is obtained via two steps. The first minimizes $\mathcal{L}_{\mathrm{reward}}$ to produce a reward function $\widehat{r}_{\widehat{\psi}}$. This is followed by a policy optimization step, which, given a language model $\widehat{\pi}_{\theta}$ parameterized by a vector $\theta$, minimizes $\mathcal{L}_{\mathrm{policy}}(\cdot|\widehat{r}_{\widehat{\psi}})$ to produce the fine-tuned language model $\widehat{\pi}_{\widehat{\theta}}$. This minimization is typically carried out via stochastic gradient descent starting with $\theta$ initialized to match the base policy $\widehat{\pi}_{\theta}=\overline{\pi}$.